kaylak:

show that f(x)=(x-2)^3+8 is one to one algebraically

Allison:

Was the other one I answered right?

kaylak:

idk im trying to finish an assignment please get me some help here

Allison:

Okay hold on.

kaylak:

ok?

Ultrilliam:

@Nnesha @sillybilly123 @Vocaloid @layla @Hero Don't mind me, just tagging our math helpers so they see this when they are on

Allison:

One to one means every x corresponds to one y, the only way I know how to do that is by evaluating every single value of x which is impossible so I don't know how to answer this question. :/

Nnesha:

If the function pass the horizontal line test then the function is 1 to 1. Like she said ^ if two different inputs give you the same outputs then the function is not 1 to 1. Algebraically you have to show that if \[f(a)=f(b) ~~~ ~~then~~a=b\] if a=b then the function is 1 to 1

Nnesha:

replace x with a replace x with b \[\large\rm f(\color{red}{a})=f(\color{blue}{b})\]\[\large\rm (\color{red}{a}-2)^3+8=(\color{blue}{b}-2)^3+8\] solve the equation for a or for b first subtract 8 both sides \[\large\rm f(\color{red}{a})=f(\color{blue}{b})\]\[\large\rm (\color{red}{a}-2)^3=(\color{blue}{b}-2)^3\] now you can take cube root both sides and simplify

Allison:

Yes! @Nnesha sorry about that, but thank you c:

Nnesha:

sorry for nothing?

Allison:

No I meant sorry for like, not knowing how to do it x.x

Nnesha:

`if two different inputs give you the same outputs then the function is not 1 to 1. Algebraically you have to show that if ` correction if two different inputs give you the same outputs then the function `is `1 to 1. Algebraically you have to show that if

Nnesha:

ohh i see. no need to say sorry especially for not knowing how to solve a question which u haven't learnd it yet

Allison:

Thank you~ :*